# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ Appendix A Solution to Remark 1
$\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}=\mathbf{H}_{i}^{T}\mathbf{H}_{i}\mathbf{z}_{i}^{t}-\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i}^{t}+\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}=-\mathbf{U}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})+c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}.$ (F.1)
@@
$=$ $\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}\big(\mathbf{z}_{i}^{t+1}+\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t}+(\mathbf{U}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})-c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t})\big)\|$
$\leq$ $\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1})\|+\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})-c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|.$ (F.2)
The above inequality follows from the triangle inequality and the non-expansiveness property of the projection operator [154]. Squaring both sides of inequality (F.2) and applying the Cauchy–Schwarz inequality further gives:
@@
$\leq$ $3\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1})\|^{2}+3c^{2}\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2}+3\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|^{2}.$ (F.3)
Now consider computing $\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1})$. Since
$\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1}):=\begin{cases}\arg\min_{\mathbf{z}}\frac{1}{2}\sum_{i\in\mathcal{N}}\|\mathbf{z}_{i}-\mathbf{z}_{i}^{t+1}\|^{2}\\ \text{subject to }\quad z\in\mathcal{X},z\in\mathcal{Y}.\end{cases}$
@@
$\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1})=\mathbf{\widetilde{W}}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})+\mathbf{z}_{i}^{t+1},\quad\text{and }\tilde{\mathbf{x}}_{i}=\mathbf{a}_{i},\ \text{if }i\in\mathcal{A}.$
@@
$\mathbf{\widetilde{W}}_{i}:=-\frac{1}{2}\cdot\mathrm{Diag}\left(\left[0,(c+1)\cdot\mathbf{1}_{N_{i}},\frac{c+1}{c}\cdot\mathbf{1}_{N_{i}},0,\mathbf{1}_{N_{i}},\mathbf{1}_{N_{i}}\right]\right)\otimes\mathbf{I}_{n}.$
Equation (F), in combination with equation (34) and the Cauchy–Schwarz inequality, leads to
$\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t+1})\|^{2}$
$=$ $\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}-\mathbf{\widetilde{W}}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}+\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1})\|^{2}$
$\leq$ $2\|(\mathbf{I}_{i}+\mathbf{\widetilde{W}}_{i})(\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1})\|^{2}+2\|\mathbf{\widetilde{W}}_{i}(\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1})\|^{2}.$ (F.7)
@@
$\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|^{2}\leq\|\mathbf{I}_{i}-\mathbf{U}_{i}\|^{2}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}.$
Substituting equations (F.7)–(F) into equation (F.3) yields
@@
$\leq$ $(6\|\mathbf{I}_{i}+\mathbf{\widetilde{W}}_{i}\|^{2}+3\|\mathbf{I}_{i}-\mathbf{U}_{i}\|^{2})\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+6\|\mathbf{\widetilde{W}}_{i}\|^{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+3c^{2}\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2}$
$=$ $\sigma_{1}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+\sigma_{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+3c^{2}(2N_{\max}+1)^{2}\|\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2},$ (F.9)
This equality holds based on the definitions of matrices $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$, $\mathbf{U}_{i}$, and $\mathbf{\widetilde{W}}_{i}$ in equations (B.1), (III), and (F), respectively, and
$\sigma_{1}$ $:=3\max\left\{\big((2(c+1)N_{\max}-1\big)^{2},(4cN_{\text{max}}-1)^{2}\right\}+6\max\left\{(\frac{1-c}{2})^{2},(\frac{1-c}{2c})^{2},(\frac{1}{2})^{2}\right\},$
@@
The definition of the function $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ in equation (IV), combined with equation (F.9) and the first part of Theorem 2, implies that $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\to 0$. Together with Lemma 2, we conclude that when the parameters satisfy equations (40)–(42), the sequence $\{(\mathbf{z}_{i},\mathbf{w}_{i},\bm{\lambda}_{i})\}_{t\geq 1}$ generated by the SP-ADMM-JCNl converges to a Karush-Kuhn-Tucker point of problem (24).
@@
Combining equation (IV) with equation (F.9) and equation (F), we obtain:
$\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\leq$ $\sum_{i\in\mathcal{N}}\left(\sigma_{1}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}+\sigma_{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+\|\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t+1}\|^{2}\right$.
@@
$\leq$ $\sum_{i\in\mathcal{N}}\big(\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}+(\sigma_{2}+2\sigma_{3}(N_{\max}+1))\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}$
@@
Here, $\sigma_{1}$ and $\sigma_{2}$ are defined in equation (F), and $\sigma_{3}=3c^{2}(2N_{\max}+1)^{2}+1$. The last inequality follows from the update step of $\bm{\lambda}_{i}^{t+1}$ and the fact that $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$. Matching the upper bound of $\mathcal{P}^{t}$ differences from equation (1) with the upper bound of $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ in equation (G), we get:
@@
where $\epsilon=\frac{\min\{c,1,\frac{\rho}{4},C_{0},C_{1},\frac{\rho}{4}-d_{\max}^{2}(\kappa_{1}+\kappa_{2})\}}{\max\{\sigma_{1},\sigma_{2}+2\sigma_{3}(1+2N_{\max}),1,\frac{2\sigma_{3}}{c^{2}}\}}$, $C_{0}:=\frac{\min\{c\kappa_{1}-6(N_{\max}+2),c(c\kappa_{1}-6(1+2c)(N_{\max}+2))\}}{\max\{3(N_{i}+2),3c(1+2c)(2+N_{\max})\}}$, and $C_{1}:=\frac{c\kappa_{2}\cdot\tilde{\tau}_{\min\quad}}{6N_{\mathrm{sum}}n(c+1)^{2}}-\frac{(2N_{max}+1)c\kappa_{1}}{2}$.
Suppose $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ reaches the lower bound $\epsilon_{1}$ for the first time at iteration $T$. Summing inequality (F) over the first $T$ iterations and combining it with equation (C.1), we have:
@@
Here, $C_{0}$ and $C_{1}$ are defined in equation (F), and $d_{\max}:=\max\{r_{i},d_{i,j},i\in\mathcal{N},j\in\mathcal{N}_{i}\}$, $N_{\max}:=\max\{N_{i},i\in\mathcal{N}\}$, $N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}$, $\tilde{\mathbf{v}}_{\min}=\min\{(c+1)^{2}N_{i}^{2}+c^{2}N_{i}+N_{i},i\in\mathcal{N}\}$.
Proof: Substituting equation (B.12) into equation (B) yields:
$\|\bm{\lambda}_{i}^{t+1}-\bm{\lambda}_{i}^{t}\|^{2}\leq\max\big\{$ $3(N_{i}+2),\,3c^{2}(2N_{\text{max}}+1)\,3c(1+2c)(2+N_{\text{max}})\big\}\cdot\big(\|\mathbf{Q}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t})-\mathbf{D}_{i}(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1})\|^{2}$
@@
1. From Algorithm 1, we have $\{\mathbf{w}^{t}\}_{t\geq 1}\subset\mathcal{B}^{\sum_{i\in\mathcal{N}}(N_{i}+1)}$, where $\mathcal{B}$ is the unit ball constraint set. Hence, $\{\mathbf{w}^{t}\}$ is bounded.

### v2

@@ A.1 Solution to Remark 2
$\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}=\mathbf{H}_{i}^{T}\mathbf{H}_{i}\mathbf{z}_{i}^{t}-\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i}^{t}+\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}=-\mathbf{U}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})-c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}.$ (F.1)
@@
$=$ $\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}\big(\tilde{\mathbf{z}}_{i}^{t+1})-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\mathbf{z}_{i}^{t}+(\mathbf{U}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})-c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t})\big)\|$
$\leq$ $\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1})\|+\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})-c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|.$ (F.2)
The above inequality follows from the triangle inequality and the non-expansiveness property of the projection operator. Squaring both sides of inequality (F.2) and applying the Cauchy–Schwarz inequality further gives:
@@
$\leq$ $3\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1})\|^{2}+3c^{2}\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2}+3\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|^{2}.$ (F.3)
Now consider computing $\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1})$. Since
$\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1}):=\begin{cases}\arg\min_{\mathbf{z}}\frac{1}{2}\sum_{i\in\mathcal{N}}\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}^{t+1}\|^{2}\\ \text{subject to }\quad z\in\mathcal{X},z\in\mathcal{Y}.\end{cases}$
@@
$\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1})=\mathbf{\widetilde{U}}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})+\tilde{\mathbf{z}}_{i}^{t+1},\quad\text{and }\tilde{\mathbf{x}}_{i}=\mathbf{a}_{i},\ \text{if }i\in\mathcal{A}.$
@@
$\mathbf{\widetilde{U}}_{i}:=-\frac{1}{2}\cdot\mathrm{Diag}\left(\left[0,(c+1)\cdot\mathbf{1}_{N_{i}},\frac{c+1}{c}\cdot\mathbf{1}_{N_{i}},0,\mathbf{1}_{N_{i}},\mathbf{1}_{N_{i}},...,0,\mathbf{1}_{N_{i}},\mathbf{1}_{N_{i}}\right]\right)\otimes\mathbf{I}_{n}.$
Equation (A.6), in combination with equation (33) and the Cauchy–Schwarz inequality, leads to
$\|\mathbf{z}_{i}^{t}-\mathrm{proj}_{\mathcal{X},\mathcal{Y}}(\tilde{\mathbf{z}}_{i}^{t+1})\|^{2}$
$=$ $\|\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{\widetilde{U}}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}+\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1})\|^{2}$
$\leq$ $2\|(\mathbf{I}_{i}+\mathbf{\widetilde{U}}_{i})(\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1})\|^{2}+2\|\mathbf{\widetilde{U}}_{i}(\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1})\|^{2}.$ (F.7)
@@
$\|(\mathbf{I}_{i}-\mathbf{U}_{i})(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|^{2}\leq\|\mathbf{I}_{i}-\mathbf{U}_{i}\|^{2}\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}.$
Substituting equations (F.7)–(A.6) into equation (F.3) yields
@@
$\leq$ $(6\|\mathbf{I}_{i}+\mathbf{\widetilde{U}}_{i}\|^{2}+3\|\mathbf{I}_{i}-\mathbf{U}_{i}\|^{2})\|\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1}\|^{2}+6\|\mathbf{\widetilde{U}}_{i}\|^{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+3c^{2}\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2}$
$=$ $\sigma_{1}\|\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1}\|^{2}+\sigma_{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+3c^{2}(2N_{\max}+1)^{2}\|\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|^{2},$ (F.9)
This equality holds based on the definitions of matrices $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$, $\mathbf{U}_{i}$, and $\mathbf{\widetilde{U}}_{i}$ in equations (A.2), (3), and (A.6), respectively, and
$\sigma_{1}$ $:=3\max\left\{4\big((c+1)N_{\max}+M_{\max}\big)^{2}-1,(4cN_{\text{max}})^{2}\right\}+6\max\left\{(\frac{1-c}{2})^{2},(\frac{1-c}{2c})^{2},(\frac{1}{2})^{2}\right\},$
@@
The definition of the function $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ in equation (4), combined with equation (F.9) and the first part of Theorem 2, implies that $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\to 0$. Together with Lemma 2, we conclude that when the parameters satisfy equations (39)–(41), the sequence $\{(\mathbf{z}_{i},\mathbf{w}_{i},\bm{\lambda}_{i})\}_{t\geq 1}$ generated by the SP-ADMM-JCNl converges to a Karush-Kuhn-Tucker point of problem (23).
@@
Combining equation (4) with equation (F.9) and equation (A.6), we obtain:
$\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\leq$ $\sum_{i\in\mathcal{N}}\left(\sigma_{1}\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}+\sigma_{2}\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}+\|\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t+1}\|^{2}\right$.
@@
$\leq$ $\sum_{i\in\mathcal{N}}\big(\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}+(\sigma_{2}+2\sigma_{3}(2N_{\max}+1))\|\mathbf{z}_{i}^{t}-\mathbf{z}_{i}^{t+1}\|^{2}$
@@
Here, $\sigma_{1}$ and $\sigma_{2}$ are defined in equation (A.6), and $\sigma_{3}=3c^{2}(2N_{\max}+1)^{2}+1$. The last inequality follows from the update step of $\bm{\lambda}_{i}^{t+1}$ and the fact that $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$. Matching the upper bound of $\mathcal{P}^{t}$ differences from equation (1) with the upper bound of $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ in equation (A.7), we get:
@@
where $\epsilon=\frac{\min\{c,1,\frac{\rho}{4},C_{0},C_{1},\frac{\rho}{4}-d_{\max}^{2}(\kappa_{1}+\kappa_{2})\}}{\max\{\sigma_{1},\sigma_{2}+2\sigma_{3}(1+2N_{\max}),1,\frac{2\sigma_{3}}{c^{2}}\}}$, $C_{0}:=\frac{\min\{c\kappa_{1}-6(N_{\max}+M_{\max}+1),c(c\kappa_{1}-6(1+c)(N_{\max}+M_{\max}+1))\}}{\max\{3(N_{i}+M_{i}+1),3c(1+c)(2+N_{\max})\}}$, and $C_{1}:=\frac{c\kappa_{2}\cdot\tilde{\tau}_{\min\quad}}{6N_{\mathrm{sum}}n(c+1)^{2}}-\frac{(2N_{max}+1)c\kappa_{1}}{2}$.
Suppose $\mathcal{G}(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ reaches the lower bound $\epsilon_{1}$ for the first time at iteration $T$. Summing inequality (A.6) over the first $T$ iterations and combining it with equation (C.1), we have:
@@
Here, $C_{0}$ and $C_{1}$ are defined in equation (A.6), and $d_{\max}:=\max\{r_{i,{\color[rgb]{0,0,0}k}},d_{i,j},i\in\mathcal{N},j\in\mathcal{N}_{i}\}$, $N_{\max}:=\max\{N_{i},i\in\mathcal{N}\}$, $N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}$, $\tilde{\mathbf{\tau}}_{\min}=\min\{(c+1)^{2}N_{i}^{2}+c^{2}N_{i}+N_{i},i\in\mathcal{N}\}$.
Proof: Substituting equation (B.12) into equation (A.2) yields:
$\|\bm{\lambda}_{i}^{t+1}-\bm{\lambda}_{i}^{t}\|^{2}\leq\max\big\{$ $3(N_{i}+M_{i}+1),\,3c^{2}(2N_{\text{max}}+1)\,3c(1+c)(1+M_{\text{max}}+N_{\text{max}})\big\}\cdot\big(\|\mathbf{Q}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t})-\mathbf{D}_{i}(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1})\|^{2}$
@@
1. From Algorithm 1, we have $\{\mathbf{w}^{t}\}_{t\geq 1}\subset\mathcal{B}^{\sum_{i\in\mathcal{N}}(N_{i}+M)}$, where $\mathcal{B}$ is the unit ball constraint set. Hence, $\{\mathbf{w}^{t}\}$ is bounded.